OpenStudy (anonymous):

Find the value of the expression "6y + 1" when y = 3

Directrix (directrix):

6y + 1 when y = 3 = 6*(3) + 1 = ?

OpenStudy (anonymous):

19

Directrix (directrix):

@shell1903 Correct.